OpenStudy (anonymous):

Judith has two closed rectangular prism recycling bins. The dimensions of the second recycle bin are one-fifth of the dimensions of the first recycle bin. Which statement is true about the surface area of the second recycle bin? It is one-twenty fifth of the surface area of the first recycle bin. It is one-thirty fifth of the surface area of the first recycle bin. It is one-fifty fifth of the surface area of the first recycle bin. It is one-forty fifth of the surface area of the first recycle bin.

OpenStudy (anonymous):

i think its the first one ?

OpenStudy (cwrw238):

if sides are in ratio 1: 5 ( first: sec on) the area are in the ratio of squares of the sides

OpenStudy (cwrw238):

right :- 5^2 = 25

OpenStudy (anonymous):

so its 1/5 ?

OpenStudy (cwrw238):

no first one says one twenty fifth ( 1 / 5^2 = 1/25)

OpenStudy (cwrw238):

u were right - its the first one